Calculate Payroll Breakin Away Company has three employees—a consultant, a computer programmer, and an administrator. The following payroll information is available for each employee: Consultant Computer Programmer Administrator Regular earnings rate $4,000 per week $60 per hour $50 per hour Overtime earnings rate* Not applicable 1.5 times hourly rate 2 times hourly rate Number of withholding allowances 2 1 2 *For hourly employees, overtime is paid for hours worked in excess of 40 hours per week. For the current pay period, the computer programmer worked 50 hours and the administrator worked 48 hours. The federal income tax withheld for all three employees, who are single, can be determined from the wage bracket withholding table in Exhibit 2. Assume further that the social security tax rate was 6.0%, the Medicare tax rate was 1.5%, and one withholding allowance is $75. Determine the gross pay and the net pay for each of the three employees for the current pay period. If required, round your answers to two decimal places. Consultant Computer Programmer Administrator Gross pay $ $ $ Net pay $ $ $
Answer #1

Ans-Determining the gross pay for each of the three employees for the current pay period as follows:

Consultant Computer Programmer Administrator
Regular earning rate $4,000 per week $60 per hour $50 per hour
Overtime earning rate NA 90 per hour (60*1.5) 100 per hour (50*2)
Number of withholding allowance 2 1 2
Hours worked 50 hours 48 hours
One withholding allowance 75
Gross Pay $4,000

$3,300

($60*40=$2,400+

$60*10*1.5=$900=

$3,300)

$2,800

($50*40=$2,000+

$50*8*2=$800=

$2,800)

Determining the net pay for each of the three employees for the current pay period as follows:

Consultant Computer Programmer Administrator
Gross Pay $4,000 $3,300 $2,800
Number of withholding allowances 2 1 2
Value of one allowance 75 75 75
Amount to be deducted 150 75 150
Amount subject to withholding $3,850 $3,225 $2,650
Less: Federal income tax withholding $939.07 $757.02 $596.02
FICA @6% (On Gross Pay) $240 $198 $168
Medicare 1.5% (On Gross Pay) $60 $49.5 $42
Net Pay $2,610.93 $2,220.48 $1,843.98

Working Note:-

Federal income tax withholding:-
Initial withholding $890.02 $356.90 $356.90
Bracket % over bracket excess

$49.05

($3,850-$3,700= $150*33%)

$400.12

($3,225-$1,796=1,429*28%)

$239.12

($2,650-$1,796=854*28%)

Total Federal income tax withholding $939.07 $757.02 $596.02